OpenStudy (anonymous):

Find both X and Y intercepts of the equation. 4x+y=5

OpenStudy (amistre64):

divide it all by 5; then divide off the top coeffs; and read off the intercepts from the bottoms

OpenStudy (amistre64):

\[4x+y=5\] \[\frac{4x}{5}+\frac{y}{5}=\frac{5}{5}\] \[\frac{x}{5/4}+\frac{y}{5}=1\]

OpenStudy (amistre64):

now when y=0; x=5/4 and when x=0; y=5